Write four solutions for each of the following equati.ons:
Ci) 2x + y = 7 (ii) me + y = 9 (iii.) x = 4y.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) Consider equation: 2x .+ y = 7 y = 7 - 2x Let x = 0, then y = 7. Solution is x = 0, y •= 7 www.cbsepdf.com LINEAR EcuATIONS IN Tw:> VARIABLES 99 Let x = 1, then y = 5. Solution is x = 1, y = 5 Let x = 2, then y = 3. Solution is x = 2, y = 3 Let x = 3, then y = 1. Solution is x = 3, y = 1. (ii) Consider equation: xx + y = 9 => y = 9 - u Let x = 0, then y = 9. Solution is x = 0, y = 9 Let x = 1, then y = 9 - Jt. Solution is x = 1, y = 9 -1t Let x = 2, then y = 9 - 21t. Solution isx = 2,y = 9 -21t Let x = 3, then y = 9 - 31t. Solution is x = 3,y = 9 -31t. (iii) Consider equation: x = 4y. Let y = 0, then x = 0. Solution is x = 0, y = 0 Let y = 1, then x = 4. Solution is x = 4, y = 1 Let y = - 1, then x = - 4. Solution is x =-4, y =- 1 Let y = 2, then x = 8. Solution is x = 8, y = 2.